# Non-overlapping, Schwarz-type Domain Decomposition Method for Physics and Equality Constrained Artificial Neural Networks

### v1

@@ Abstract
We introduce a non-overlapping, Schwarz-type domain decomposition method employing a generalized interface condition, tailored for physics-informed machine learning of partial differential equations (PDEs) in both forward and inverse scenarios. Our method utilizes physics and equality constrained artificial neural networks (PECANN) in each subdomain. Diverging from the original PECANN method, which uses initial and boundary conditions to constrain the PDEs alone, our method jointly employs both the boundary conditions and PDEs to constrain a specially formulated generalized interface loss function for each subdomain. This modification enhances the learning of subdomain-specific interface parameters, while delaying information exchange between neighboring subdomains, and thereby significantly reduces communication overhead. By utilizing an augmented Lagrangian method with a conditionally adaptive update strategy, the constrained optimization problem in each subdomain is transformed into a dual unconstrained problem. This approach enables neural network training without the need for ad-hoc tuning of model parameters. We demonstrate the generalization ability and robust parallel performance of our method across a range of forward and inverse problems, with solid parallel scaling performance up to 32 processes using the Message Passing Interface model. A key strength of our approach is its capability to solve both Laplace’s and Helmholtz equations with multi-scale solutions within a unified framework, highlighting its broad applicability and efficiency.
@@
, Augmented Lagrangian method , constrained optimization , domain decomposition , physics-informed neural networks
journal:
@@
[inst1]organization=Department of Mechanical Engineering and Materials Science, University of Pittsburgh,city=Pittsburgh, postcode=15261, state=PA, country=USA
@@ 1 Introduction
Jagtap et al. [19] decomposed a spatial domain into smaller domains and used the baseline PINN method to learn the solution of a PDEs on the whole domain. A separate neural network was adopted in each subdomain and flux continuity across subdomain interfaces were enforced in strong form. The average value of the solution between two subdomains sharing an interface was also enforced as an additional condition. Since, the neural network models associated for subdomains exchange information at each epoch, makes it not strictly a Schwarz-type DDM. In the spirit of the baseline PINNs, loss terms arising from the flux continuity across subdomain interfaces are lumped into a single composite objective function with tunable weights. In a followup work, Jagtap and Karniadakis [20] extended the work presented in [19] to include the time domain. Furthermore, in this followup work, the interface conditions were simplified to make the method applicable to PDEs that may not represent conservation laws. Hu et al. [21] explored the conditions and mechanisms by which Extended PINN (XPINN) with domain decomposition enhances generalization compared to PINN without domain decomposition. Their analysis indicated that domain decomposition can improve predictions by breaking the overall problem into simpler, more manageable subproblems. However, they also noted that this strategy reaches a limit as the availability of training points decreases with an increasing number of subdomains. A parallel implementation of these works is presented in [22] showing decent scalability and speedup.
@@
Clearly, domain decomposition in the context of scientific machine learning or physics-informed neural networks is a growing area of focus, because it enables neural networks to tackle larger and complex problems or reduce training times substantially. Additionally, empirical evidence shows that training separate neural networks on smaller domains is much more feasible and likely to converge than training a single neural network on a large domain with many points [21], a feature that we confirm in the present study as well. In what follows, we present the theory behind the optimized Schwarz methods [11] as it inspires our proposed approach. We then propose a non-overlapping, Schwarz-type domain decomposition method with an interface condition with learnable, subdomain-specific parameters. We apply the resulting distributed learning method to solve forward and inverse PDE problems on multiple processors and assess its parallel performance up to 32 processes.
@@ 2 Optimized Schwarz method
where $\bm{n_{1}}$ and $\bm{n_{2}}$ are the outward normal directions on the subdomain boundaries of $\Omega_{1}$ and $\Omega_{2}$, respectively. Note that $\bm{n_{1}}=-\bm{n_{2}}$. $\Gamma_{1}$ and $\Gamma_{2}$ represents the subdomain interfaces corresponding to $\Omega_{1}$ and $\Omega_{2}$, respectively. In the case of a non-overlapping domain decomposition $\Gamma_{1}$ and $\Gamma_{2}$ are identical. $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are operators act along the interfaces $\Gamma_{1}$ and $\Gamma_{2}$, respectively. $\beta_{1}$ and $\beta_{2}$ are real valued functions. With $\beta_{1}=\beta_{2}=0$ and $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ being identity operators, the original Schwarz method is recovered. As a remedy to the drawbacks of classical Schwarz methods, Lions [26] proposed to replace Dirichlet interface conditions with Robin interface conditions with a tunable parameter $\alpha$. In the above interface formulation, we see that with $\beta_{1}=\beta_{2}=1$ and $\mathcal{A}_{1}=\mathcal{A}_{2}=\alpha$, where $\alpha>0.0$, we recover the Robin interface conditions proposed by Lions [26].
@@ 3 Technical Formulation
Our primary objective is to determine the solutions of both forward and inverse partial differential equation (PDE) problems in a distributed manner using neural networks. To achieve this, we pursue a constrained optimization formalism and develop a Schwarz-type DDM and implement it using the Message Passing Interface (MPI) standard. Specifically, our proposed method extends the constrained optimization formulation of the Physics and Equality Constrained Artificial Neural Networks (PECANN) framework [25] to learn the interface transmission conditions arising from domain decomposition. A distinctive feature of our approach is the simultaneous learning of the interface transmission conditions and the solution of the PDE within each subdomain. The efficacy of this approach is demonstrated by its ability to solve multi-scale Poisson’s and Helmholtz equations, as well as inverse PDE problems.
@@ 3.1 PECANN: Physics and Equality Constrained Artificial Neural Networks
$\mathcal{M}(\textbf{x};\theta)$ $=\mathcal{D}(\textbf{x};\theta),\quad$ $\forall\textbf{x}\in M_{k}\subset\Omega_{k},$
where $\alpha,\beta,\gamma$ are subdomain specific parameters of the interface condition.
@@
It is worthwhile to note a critical aspect of the formulation presented in Eq. 8. The objective function $\mathcal{J}$ incorporates an interface transmission condition, while the constraint function $\mathcal{C}_{j}$, $j\in\{B,P,M\}$, encompasses boundary and/or data constraints, along with physics constraints that represent the governing equation at hand. This feature sets our approach apart from the original PECANN [25], PINN [33], and its distributed variant XPINN [20, 21] frameworks. Notably, we treat the governing PDE as a constraint, rather than the subdomain interface conditions. We find this approach to be much more robust than using interface conditions as constraints, particularly because interface predictions can be inaccurate during the initial stages of training when the interface parameters are still undetermined.
@@ 3.2 Formulation of the Interface Loss Function
To analyze the optimality conditions of the objective function $\mathcal{J}$ with respect to the interface parameters in conditions given in Eq. 3, we consider the derivative of the interface loss at the $i$th collocation point $\textbf{x}_{i}$ on a subdomain interface. In vector form, the interface operator is rewritten as:
$\mathcal{I}_{i}=\textbf{q}^{T}\bm{\mathcal{O}}_{i},$ (9)
@@
where the interface parameters are denoted by $\textbf{q}=[\alpha,\beta,\gamma]^{T}$ and the interface operators by $\bm{\mathcal{O}}_{i}=[\mathcal{D}_{i},\mathcal{N}_{i},\mathcal{T}_{i}]^{T}$. The gradient of the loss function with respect to the interface parameters is given by:
$\frac{\partial\mathcal{J}_{i}}{\partial\textbf{q}}=2\mathcal{I}_{i}\bm{\mathcal{O}}_{i}.$ (10)
@@
At optimality, the gradient of the loss function must vanish, so we expect either $\mathcal{I}_{i}$ or $\bm{\mathcal{O}}_{i}$ for the $i$th interface to be zero. The presence of $\mathcal{I}_{i}$ inherently couples the Dirichlet, Neumann, and continuity of tangential derivative conditions. This coupling implies that the elements of $\mathbf{q}$ are not truly independent, making it difficult to reach optimality in practice.
Expanding the interface loss, we have:
$\mathcal{J}_{i}=\mathcal{I}_{i}^{2}=(\textbf{q}^{T}\bm{\mathcal{O}}_{i})^{2}=\sum_{j}(q^{j}\mathcal{O}_{i}^{j})^{2}+2\sum_{j\neq m}q^{j}q^{m}\mathcal{O}_{i}^{j}\mathcal{O}_{i}^{m},$ (11)
which we refer to as the full interface loss in the appendix. The cross-product terms (i.e. $q^{j}q^{m}\mathcal{O}_{i}^{j}\mathcal{O}_{i}^{m}$) in Eq. 11 is the root cause of this problematic dependency, which arises from the implementation of the MSE metric. Although the MSE metric has been deemed efficient in PINNs, it introduces convergence issues when applied to the generalized interface transmission condition for optimal satisfaction of each interface continuity. To address this, we consider the following two alternatives to break the dependency and enhance convergence.
@@
The first approach involves using an absolute distance function. Although less efficient, it avoids the cross-product terms, thereby preventing the coupling of interface parameters. It is expressed as:
$\mathcal{J}_{i}=|\mathcal{I}_{i}|=\sum_{j}q^{j}|\mathcal{O}_{i}^{j}|,$ (12)
@@
which we refer to as the absolute interface loss. While not our proposed solution, this formulation illustrates an alternative means of representing the generalized interface condition in the loss function.
To address the paradoxical situation between dependency and efficiency induced by the MSE metric, we propose the following alternative implementation: neglecting the cross-product terms in Eq. 11 to approximate the interface loss function as follows:
$\mathcal{J}_{i}$ $=\mathcal{I}_{i}^{2}\approx\sum_{j}(q^{j}\mathcal{O}_{i}^{j})^{2}.$ (13)
@@
In the appendix, we will refer to it as approximate interface loss. With this approximation, the optimality condition can now be expressed in a clear manner:
$\frac{\partial\mathcal{J}_{i}}{\partial\textbf{q}}=2\textbf{q}\odot\bm{\mathcal{O}}_{i}\odot\bm{\mathcal{O}}_{i},$ (14)
@@
where $\odot$ denotes an element-wise product. Again, at optimality, we expect each element of $\bm{\mathcal{O}}_{i}$ to vanish. Therefore, each element of q do indeed become independent parameters, which aligns with our aim. The independent adjustability of the interface parameters facilitates proper adaptation to the distinct physical scales inherent in their corresponding operators. We find this adaptability to be key for effectively tackling both Poisson’s and Helmholtz equations using the same framework of PECANN. Note that the objective function will be unique to the type of PDE problem because the interface loss is constrained by the PDE. Moreover, when $\mathcal{O}_{i}^{j}$ is small or converges to small values, increasing the corresponding $q^{j}$ helps maintain a large gradient, thus enhancing the convergence rate. This adjustment leverages the efficiency of the gradient-based optimizer, improving the overall performance of the solution process. In other words, these parameters serve as multipliers for the Dirichlet, Neumann and tangential derivative continuity operators, respectively. However, it is crucial to emphasize that the method by which these parameters are predicted in our approach differs fundamentally from their determination in optimized Schwarz methods. To further analyze the convergence characteristics of interface parameters as multipliers, we investigate the ratios $\dfrac{\beta}{\alpha}$ and $\dfrac{\gamma}{\alpha}$ for the $k$th subdomain. These ratios are expected to asymptotically converge to constant values towards optimality.
@@ 3.3 Conditionally Adaptive Augmented Lagrangian Method
$\max_{\textbf{q}>1}\min_{\theta}\mathcal{J}(\theta,\textbf{q};\textbf{x})\approx\frac{1}{N_{\Gamma_{k}}}\sum_{i=1}^{N_{\Gamma_{k}}}\sum_{j}\left(q^{j}\mathcal{O}^{j}(\theta;\textbf{x}_{i})\right)^{2}$ on $\leavevmode\nobreak\ \text{on }\Gamma_{k}.$ (15)
@@
$\max_{\textbf{q}>1,\bm{\lambda}}\min_{\theta}\mathcal{L}(\theta,\bm{\lambda},\textbf{q};\textbf{x},\bm{\mu})=\mathcal{J}(\theta,\mathbf{q};\textbf{x})+\bm{\lambda}^{T}\bm{\mathcal{C}}(\theta;\textbf{x})+\frac{1}{2}\bm{\mu}^{T}\big{(}\bm{\mathcal{C}}(\theta;\textbf{x})\odot\bm{\mathcal{C}}(\theta;\textbf{x})\big{)},$ (16)
where $\bm{\lambda}$ and $\bm{\mu}$ are the vectors of Lagrange multipliers and penalty parameters corresponding to the vector of constraints $\bm{\mathcal{C}}=[\mathcal{C}_{B},\mathcal{C}_{P},\mathcal{C}_{D}]^{T}$, respectively. Note that unlike in the conventional ALM, we employ a vector containing of independent penalty parameters to address the different characteristics of the constraints.
The minimization of Eq. 16 can be performed using a variant of gradient descent type optimizer for Lagrange multipliers, inspired by the adaptive augmented Lagrangian method in Basir and Senocak [36]. This process is outlined in Algorithm 1. In each outer iteration of the algorithm, a specific unconstrained optimization problem, given $\bm{\lambda}^{p-1}$ and $\bm{\mu}^{p-1}$, is solved to satisfy a convergence criterion over a limited number of inner epochs $N$ through training. The solution is considered converged if the current augmented Lagrangian loss function $\mathcal{L}^{p}$ exceeds a fraction $\omega$ (default value 0.999) of the previous loss $\mathcal{L}^{p-1}$. Upon convergence, updates to the Lagrange multipliers and penalty parameters are made, and the process proceeds to the next unconstrained optimization problem with updated $\bm{\lambda}^{p}$ and $\bm{\mu}^{p}$. For domain decomposition, an additional condition requires that the optimization progress meets a minimum epoch threshold, $epoch_{min}$, to prevent excessive communication between the adjacent subdomains.
@@
3for $n=1\leavevmode\nobreak\ \textnormal{{to}}\leavevmode\nobreak\ N$ do
4       for $p=1\leavevmode\nobreak\ \textnormal{{to}}\leavevmode\nobreak\ P$ do
             $\theta_{k}^{p}\leftarrow\text{Train}(\theta_{k}^{p-1})$                                  /* primal update */
@@
5             if $\mathcal{L}^{p}>=\omega\mathcal{L}^{p-1}$ then
@@
6                   if DDM == ON AND $p<epoch_{min}$ then
7                         continue
8                   else
@@
It is crucial to highlight the robustness of the PECANN framework in formulating and solving both forward and inverse PDE problems through a constrained optimization formalism. This stands in stark contrast to the unconstrained optimization formalism adopted in the PINN approach. PECANN strictly adheres to a constrained optimization formalism, from which an equivalent dual unconstrained optimization problem is derived using the augmented Lagrangian method [34, 35]. This distinctive approach enable PECANN to integrate diverse constraints in the learning process systematically, avoiding any heuristic methods to balance terms in a composite objective function that arises due to an unconstrained optimization formalism.

### v2

@@ Abstract
We present a non-overlapping, Schwarz-type domain decomposition method with a generalized interface condition, designed for physics-informed machine learning of partial differential equations (PDEs) in both forward and inverse contexts. Our approach employs physics and equality-constrained artificial neural networks (PECANN) within each subdomain. Unlike the original PECANN method, which relies solely on initial and boundary conditions to constrain PDEs, our method uses both boundary conditions and the governing PDE to constrain a unique interface loss function for each subdomain. This modification improves the learning of subdomain-specific interface parameters while reducing communication overhead by delaying information exchange between neighboring subdomains. To address the constrained optimization in each subdomain, we apply an augmented Lagrangian method with a conditionally adaptive update strategy, transforming the problem into an unconstrained dual optimization. A distinct advantage of our domain decomposition method is its ability to learn solutions to both Poisson’s and Helmholtz equations, even in cases with high-wavenumber and complex-valued solutions. Through numerical experiments with up to 64 subdomains, we demonstrate that our method consistently generalizes well as the number of subdomains increases.
@@
, Augmented Lagrangian method , , domain decomposition , Helmholtz equation , physics-informed neural networks , Poisson’s equation
@@
[inst1]organization=Department of Mechanical Engineering and Materials Science,
University of Pittsburgh,city=Pittsburgh, postcode=15261, state=PA, country=USA
@@ 1 Introduction
Jagtap et al. [19] decomposed a spatial domain into smaller domains and used the baseline PINN method to learn the solution of a PDEs on the whole domain. A separate neural network was adopted in each subdomain and flux continuity across subdomain interfaces were enforced in strong form. The average value of the solution between two subdomains sharing an interface was also enforced as an additional condition. Since, the neural network models associated for subdomains exchange information at each epoch, makes it not strictly a Schwarz-type DDM. In the spirit of the baseline PINNs, loss terms arising from the flux continuity across subdomain interfaces are lumped into a single composite objective function with tunable weights. In a followup work, Jagtap and Karniadakis [20] extended the work presented in [19] to include the time domain. Furthermore, in this followup work, the interface conditions were simplified to make the method applicable to PDEs that may not represent conservation laws. However, challenges related to the interface loss term in XPINN were acknowledged in a recent work and a trainable gate network with a flexible domain decomposition strategy was proposed [21]. Hu et al. [22] explored the conditions and mechanisms by which Extended PINN (XPINN) with domain decomposition enhances generalization compared to PINN without domain decomposition. Their analysis indicated that domain decomposition can improve predictions by breaking the overall problem into simpler, more manageable subproblems. However, they also noted that this strategy reaches a limit as the availability of training points decreases with an increasing number of subdomains. A parallel implementation of XPINN was presented in [23] showing decent scalability and speedup.
@@
Clearly, domain decomposition in the context of scientific machine learning or physics-informed neural networks is a growing area of focus, because it enables neural networks to tackle larger and complex problems or reduce training times substantially. Additionally, empirical evidence shows that training separate neural networks on smaller domains is much more feasible and likely to converge than training a single neural network on a large domain with many points [22], a feature that we confirm in the present study as well. In what follows, we present the theory behind the optimized Schwarz methods [11] as it inspires our proposed approach. We then propose a non-overlapping, Schwarz-type domain decomposition method with an interface condition with learnable, subdomain-specific parameters. We apply the resulting distributed learning method to forward and inverse PDE problems. Notably, our examples include Poisson’s and Helmholtz equations, with high wavenumber and complex-valued solutions.
@@ 2 Optimized Schwarz method
where $\Delta$ is the Laplacian operator applied to the function $u$, and $s_{1,2}$ are given source terms. $\bm{n_{1}}$ and $\bm{n_{2}}$ are the outward normal directions on the subdomain boundaries of $\Omega_{1}$ and $\Omega_{2}$, respectively. Note that $\bm{n_{1}}=-\bm{n_{2}}$. $\Gamma_{1}$ and $\Gamma_{2}$ represents the subdomain interfaces corresponding to $\Omega_{1}$ and $\Omega_{2}$, respectively. In the case of a non-overlapping domain decomposition $\Gamma_{1}$ and $\Gamma_{2}$ are identical. $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are operators act along the interfaces $\Gamma_{1}$ and $\Gamma_{2}$, respectively. $\beta_{1}$ and $\beta_{2}$ are real valued functions. With $\beta_{1}=\beta_{2}=0$ and $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ being identity operators, the original Schwarz method is recovered. As a remedy to the drawbacks of classical Schwarz methods, Lions [27] proposed to replace Dirichlet interface conditions with Robin interface conditions with a tunable parameter $\alpha$. In the above interface formulation, we see that with $\beta_{1}=\beta_{2}=1$ and $\mathcal{A}_{1}=\mathcal{A}_{2}=\alpha$, where $\alpha>0.0$, we recover the Robin interface conditions proposed by Lions [27].
@@ 3 Technical Formulation
Our primary objective is to determine the solutions of both forward and inverse partial differential equation (PDE) problems in a distributed manner using neural networks. To achieve this, we pursue a constrained optimization formalism and develop a non-overlapping Schwarz-type DDM and implement it using the Message Passing Interface (MPI) standard. Specifically, our proposed method extends the constrained optimization formulation of the Physics and Equality Constrained Artificial Neural Networks (PECANN) framework [26] to a generalized interface loss term arising from domain decomposition. A distinctive feature of our approach is the simultaneous learning of the parameters of the interface conditions and the solution of the PDE within each subdomain. We demonstrate the distinct advantege of this by its ability to learn the solutions of both Poisson’s and Helmholtz equations, with high wavenumber and complex-valued solutions. We also apply our method to inverse PDE problems.
@@ 3.1 PECANN: Physics and Equality Constrained Artificial Neural Networks
$\mathcal{M}(\textbf{x};\theta)$ $=\mathcal{D}(\textbf{x};\theta),\quad$ $\forall\textbf{x}\in M_{k}\subset\Omega_{k}.$
@@
It is worthwhile to note a critical aspect of the formulation presented in Eq. 8. The objective function $\mathcal{J}$ incorporates an interface transmission condition, while the constraint function $\mathcal{C}_{j}$, $j\in\{B,P,M\}$, encompasses boundary and/or data constraints, along with physics constraints that represent the governing equation at hand. This feature sets our approach apart from the original PECANN [26], PINN [34] and its distributed variant XPINN [20, 22] frameworks. Notably, we treat the governing PDE as a constraint, rather than the subdomain interface conditions. We find this approach to be much more robust than using interface conditions as constraints, particularly because interface predictions can be inaccurate during the initial stages of training when the interface parameters are still undetermined.
@@ 3.2 Formulation of the Interface Loss Function
To analyze the optimality conditions of the objective function $\mathcal{J}$ with respect to the interface parameters in conditions given in Eq. 3, we consider the derivative of the interface loss at the $i$th collocation point $\textbf{x}_{i}$ on a subdomain interface. In summation form, the interface operator is rewritten as:
$\mathcal{I}_{i}=\sum_{j=1}^{3}q^{j}\mathcal{O}_{i}^{j},$ (9)
@@
where $q^{j}$ is the $j$th interface parameter in $\textbf{q}=[\alpha,\beta,\gamma]^{T}$ and $\mathcal{O}_{i}^{j}$ is the corresponding interface operator in $\bm{\mathcal{O}}_{i}=[\mathcal{D}_{i},\mathcal{N}_{i},\mathcal{T}_{i}]^{T}$. Since we adopt the MSE as a metric, the loss function will be $\mathcal{J}_{i}=\mathcal{I}_{i}^{2}$. The gradient of the loss function with respect to the $j$th interface parameter is given by:
$\frac{\partial\mathcal{J}_{i}}{\partial q^{j}}=2\mathcal{O}_{i}^{j}\mathcal{I}_{i}=2\mathcal{O}_{i}^{j}\sum_{j=1}^{3}q^{j}\mathcal{O}_{i}^{j}.$ (10)
@@
At optimality, we expect the above gradient of the loss function to vanish, meaning either $\mathcal{I}_{i}$ or $\mathcal{O}_{i}^{j}$ for the $i$th interface must be zero. The presence of $\mathcal{I}_{i}$ in Eq. 10 naturally couples the Dirichlet, Neumann, and continuity of tangential derivative conditions with interface parameters q. The presence of $\mathcal{I}_{i}$ inherently couples the Dirichlet, Neumann, and continuity of tangential derivative conditions. This coupling implies that the elements of $\mathbf{q}$ are not truly independent, hindering the attainment of optimality in practice. To address this issue, we introduce an approximation to the interface loss function, as detailed below.
Let us start with expanding the interface loss as follows:
$\mathcal{J}_{i}=\sum_{j=1}^{3}(q^{j}\mathcal{O}_{i}^{j})^{2}+2\sum_{j\neq m}q^{j}q^{m}\mathcal{O}_{i}^{j}\mathcal{O}_{i}^{m},$ (11)
which we will refer to as the full interface loss. When taking the gradient of the loss function $\mathcal{J}_{i}$ with respect to interface parameters q, the cross-product terms (i.e. $q^{j}q^{m}\mathcal{O}_{i}^{j}\mathcal{O}_{i}^{m}$) in Eq. 11 are responsible for this problematic coupling issue. To address this issue, we consider the following two alternatives to break the dependency and enhance convergence.
@@
Our first approach involves using an absolute distance function metric to formulate a loss function. Although less efficient, the absolute distance function metric does not lead to the cross-product terms, thereby preventing the coupling of interface parameters. The loss function is expressed as:
$\mathcal{J}_{i}=|\mathcal{I}_{i}|=\sum_{j=1}^{3}q^{j}|\mathcal{O}_{i}^{j}|,$ (12)
@@
which we refer to as the absolute interface loss. While this is not our proposed solution, the formulation illustrates an alternative means of representing the generalized interface condition as a loss term for the optimization problem.
For our second approach, we propose the following approximation to Eq. 11:
$\mathcal{J}_{i}$ $=\mathcal{I}_{i}^{2}\approx\sum_{j=1}^{3}(q^{j}\mathcal{O}_{i}^{j})^{2}.$ (13)
@@
In arriving at the above approximation, we neglected the cross-product terms in Eq. 11. We will refer to Eq. 13 as approximate interface loss. With this approximation to the loss function, the optimality condition can now be expressed as
$\frac{\partial\mathcal{J}_{i}}{\partial q^{j}}$ $=2q^{j}(\mathcal{O}_{i}^{j})^{2}.$ (14)
Again, at optimality, we expect each element of $\bm{\mathcal{O}}_{i}$ to vanish. Therefore, each element of q does indeed become independent parameters, which aligns with our aim. We should also note that the independent learning of the interface parameters facilitates proper adaptation to the distinct physical scales or dimensions inherent in their corresponding operators.
@@
We find the use Eq. 13 as a generalized interface loss term and the adaptability of its subdomain specific parameters to be key for effectively tackling both Poisson’s and Helmholtz equations using the same PECANN framework. We note that the resulting objective function will be unique to the type of PDE problem because the interface loss is constrained by the target PDE and its boundary conditions. Moreover, when $\mathcal{O}_{i}^{j}$ is small or converging toward to small values, a larger $q^{j}$ helps maintain a large gradient, thus enhancing the convergence rate. This adjustment leverages the efficiency of the gradient-based optimizer, improving the overall performance of the solution process. In other words, these parameters serve as multipliers for the Dirichlet, Neumann and tangential derivative continuity operators, respectively. However, it is crucial to emphasize that the method by which these parameters are predicted in our approach differs fundamentally from their determination in optimized Schwarz methods.
To further examine the convergence behavior of the interface parameters as multipliers, we analyze the ratios $\dfrac{\beta}{\alpha}$ and $\dfrac{\gamma}{\alpha}$ for the $k$th subdomain. These ratios are expected to asymptotically stabilize to constant values upon the method’s convergence. This expectation follows from Eq. 3 and Eq. 4, where the interface conditions should be satisfied at convergence. Consequently, we can treat one of the parameters as a multiplier; normalizing both sides by this parameter yields two independent ratios that the network must learn.
@@ 3.3 Conditionally Adaptive Augmented Lagrangian Method
$\min_{\theta}\max_{\textbf{q}>1}\mathcal{J}(\theta,\textbf{q};\textbf{x})\approx\frac{1}{N_{\Gamma_{k}}}\sum_{i=1}^{N_{\Gamma_{k}}}\sum_{j=1}^{3}\left(q^{j}\mathcal{O}^{j}(\theta;\textbf{x}_{i})\right)^{2}$ on $\leavevmode\nobreak\ \text{on }\Gamma_{k}.$ (15)
@@
$\min_{\theta}\max_{\textbf{q}>1,\bm{\lambda}}\mathcal{L}(\theta,\bm{\lambda},\textbf{q};\textbf{x},\bm{\mu})=\mathcal{J}(\theta,\mathbf{q};\textbf{x})+\bm{\lambda}^{T}\bm{\mathcal{C}}(\theta;\textbf{x})+\frac{1}{2}\bm{\mu}^{T}\big{(}\bm{\mathcal{C}}(\theta;\textbf{x})\odot\bm{\mathcal{C}}(\theta;\textbf{x})\big{)},$ (16)
where $\odot$ denotes an element-wise product, and $\bm{\lambda}$ and $\bm{\mu}$ are the vectors of Lagrange multipliers and penalty parameters corresponding to the vector of constraints $\bm{\mathcal{C}}=[\mathcal{C}_{B},\mathcal{C}_{P},\mathcal{C}_{D}]^{T}$, respectively. Note that unlike in the conventional ALM, we employ a vector containing of independent penalty parameters to address the different characteristics of the constraints.
The minimization of Eq. 16 can be performed using a variant of gradient descent type optimizer for Lagrange multipliers, inspired by the adaptive augmented Lagrangian method in Basir and Senocak [37]. This process is outlined in Algorithm 1. In each outer iteration of the algorithm, a specific unconstrained optimization problem, given $\bm{\lambda}^{p-1}$ and $\bm{\mu}^{p-1}$, is solved to satisfy a convergence criterion over a limited number of inner epochs $P$ through training. The solution is considered converged if the current augmented Lagrangian loss function $\mathcal{L}^{p}$ exceeds a fraction $\omega$ (default value 0.999) of the previous loss $\mathcal{L}^{p-1}$. Upon convergence, updates to the Lagrange multipliers and penalty parameters are made, and the process proceeds to the next unconstrained optimization problem with updated $\bm{\lambda}^{p}$ and $\bm{\mu}^{p}$. For domain decomposition, an additional condition requires that the optimization progress meets a minimum epoch threshold, $epoch_{min}$, to prevent excessive communication between the adjacent subdomains.
@@
3for $n\leftarrow 1\leavevmode\nobreak\ \textnormal{{to}}\leavevmode\nobreak\ N$ do
       $\mathcal{L}^{0}\leftarrow\infty$                       /* Initializing augmented Lagrangian loss */
4       for $p\leftarrow 1\leavevmode\nobreak\ \textnormal{{to}}\leavevmode\nobreak\ P$ do
             $\theta_{k}^{p}\leftarrow\text{Update}(\theta_{k}^{p-1})$                                  /* primal update */
             $\mathcal{L}^{p}\leftarrow$ Eq. 16 with $\theta_{k}^{p},\bm{\lambda}^{p-1},\bm{\mu}^{p-1},\cdots$       /* augmented loss update */
@@
5             if $\mathcal{L}^{p}>=\omega\mathcal{L}^{p-1}$ OR $p==P$ then
@@
6                   if $p\geq epoch_{min}$ then
@@
It is crucial to highlight the robustness of the PECANN framework in formulating and solving both forward and inverse PDE problems through a constrained optimization formalism. This stands in stark contrast to the unconstrained optimization formalism adopted in the PINN approach. PECANN strictly adheres to a constrained optimization formalism, from which an equivalent dual unconstrained optimization problem is derived using the augmented Lagrangian method [35, 36]. This distinctive approach enables PECANN to systematically integrate diverse constraints into the learning process, eliminating the need for heuristic methods to balance terms within a composite objective function, as would be required in an unconstrained optimization framework.